40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 194 of 344  Not logged in ELOG logo
ID Date Author Type Category Subject
  7570   Wed Oct 17 19:35:58 2012 KojiUpdateComputersRe: Lots of new White :(

Solved. The power code of c1iscaux was loose.
Has anyone worked around the back side of 1Y3?


I looked into the problem. I went around the channel lists for each slow machines and found the variables are supported by c1iscaux

controls@pianosa:/cvs/cds/caltech/target/c1iscaux 0$ cd /cvs/cds/caltech/target/c1iscaux
controls@pianosa:/cvs/cds/caltech/target/c1iscaux 0$ grep C1:IF *
C1IFO_STATE.db:grecord(ai,"C1:IFO-STATE")

It seemed that the machine was not responding to ping. I went to 1Y3 and found the crate was off. Actually this is not correct.
The key was on but the power was off. I looked at the back and found the power code was loose from its inlet.
Once the code was pushed in and the crate was keyed, the white boxes got back online.

Just in case I burtrestored these slow channels by the snapshot at 6:07am on Sunday.

  7569   Wed Oct 17 18:41:27 2012 JenneUpdateCamerasCamera looking at ETMY baffle

The camera titled "watec_mobile" is looking at the front of the black glass baffle (i.e. the side facing the ITM) on the ETMY table.  This required (for my quick hacky solution) removing the regular ETMYF camera.  Steve has a genius plan (I think) so that we can have both at the same time.  Anyhow, eventually we'll move the black glass back, so we'll be back to needing just one camera.

After dinner, I'll try aligning the Yarm.

  7568   Wed Oct 17 11:15:05 2012 DenUpdateSUS PITCH damping needed

Quote:

 

 Pitch damping solution needed! It should be in the machine shop already.

I think we can put ø2mm × 10mm long magnetic material inside 4 holes with actuation magnets. Then magnetic field on the other side of the mirror will be close to one produced by actuation magnet. Magnetic cylinder center of inertia will be in the vertical plane where mirror's center of inertia is. So this should not change alignment significantly. Eddy current dumping will be applied to the end of the magnetic cylinder opposite to the magnet using aluminium disks, we have them in the clean room.

pitch_damping.png

  7567   Wed Oct 17 09:01:37 2012 SteveUpdatePEMants on the PSL table

Quote:

We observed one or two ants climbing over PMC optics without booties and safety glasses.

The floor was mopped with strong Bayer Home Pest Control solution in the Vertex area.

Do not work inside the 40m lab if you are sensitive to chemicals!

 

 The floor is mopped again with strong PEST CONTROL SOLUTION in water in the Vertex area.

Do not plane to work in the IFO-room till noon if you are sensitive to chemicals!

  7566   Wed Oct 17 08:30:49 2012 SteveUpdateSUS PTICH damping needed

Quote:

Quote:

  [Koji, Steve, Den]

TT alignment is fine, yaw damping is satisfactory, pitch damping is slow. We might want to add magnets to the mirror and attach blades to the frame for pitch edge current damping.

We are moving towards electronics testing.

Atm1,  TT 1.5" high adaptor base will be back from the shop in 10 days.

Atm2,  There is no PITCH damping, YAW edie current damping works well at 0.5 mm gap

Atm3,  Adjustable Al -disc that contains a small magnet is purely designed.

We have to come up with a solution to have damping in PITCH

 Pitch damping solution needed! It should be in the machine shop already.

  7565   Wed Oct 17 08:05:10 2012 DenUpdateGeneraldo not leave op-table open

Quote:

ETMX optical table was left open. Burned toast award goes to ......?

 me

  7564   Wed Oct 17 08:04:31 2012 SteveUpdateVACRoughing pumps on for oil change

PR1, PR2 and RP3 turned on for warming up for oil change. Oil changed with 3.2L of  MVT-19 fluid in each. This substitute for HE-175 will be used at the next oil change - it has 1E-6 Torr vapor pressure.

To finish this job tomorrow: 1, check oil creeping upstream  2, change air filter of air purge if pressure drops <350 mTorr  3, measure venting time of pump

  7563   Wed Oct 17 08:01:54 2012 SteveUpdateGeneraldo not leave op-table open

ETMX optical table was left open. Burned toast award goes to ......?

  7562   Tue Oct 16 21:57:47 2012 DenUpdatePEMaccelerometers

All accelerometers are now at the table behind 1X4, cables are near readout box.

  7561   Tue Oct 16 20:40:06 2012 DenUpdateIOOc1lsc DAC0 now connected to tip-tilt SOS DW boards

  Full cable path from coil driver to osem input is now ready. I've tested Ch1-4 of the left AI and left coil driver. 15 pin outputs and monitors show voltage that we expect. I've checked voltage on the other side of the cable in the clean room, it is correct. We are ready to test the coils. We need to bake osem cables asap. Hopefully, Bob will start this job tomorrow.

DSC_4749.JPG DSC_4755.JPG

  7560   Tue Oct 16 17:13:23 2012 CzarinaSummaryGeneralvent stuff - 4 paths

I see 4+ possible paths for us to take, in terms of a possible vent in the next few weeks:

No Vent - Just do FPPRMI, using AS55

Mini Vent - Fix REFL path, nothing else.  ~1 day at atmosphere

Medium Vent - Fix REFL path, swap G&H mirrors for LaserOptik mirrors (so also resuspend passive TTs, maybe add pitch adjustment option). ~1 week or so at atmosphere - do this rather than Mini if Jan's Finesse calc says the G&H mirrors are too rough

Mega Vent - Fix all the things, do all the things.  Long time at atmosphere

The "+" is to take into account all the possible variations on "medium vent".  The No, Mini and Medium options assume we'll do the Mega option later, just not immediately.

  7559   Tue Oct 16 17:06:47 2012 SteveUpdateSAFETYsafety training

Ayaka Shoda, visiting graduate student received basic 40m specific safety training today

Attachment 1: IMG_1735.JPG
IMG_1735.JPG
  7558   Tue Oct 16 16:46:20 2012 Vent Czar and CzarinaSummaryGeneralvent stuff

We're discussing the plan for the next vent.  Now that we have all the tip-tilt stuff in hand, we should get ready for the big TT installation vent ASAP.  The question is what remains to be done, and when can we be ready

Stuff that needs doing:

* characterize TTs, check electrical connection of quadrupus cables, bake cables and base plates - a week or so (not including baking)

* phase map of LaserOptik mirrors (decide on needed resolution) - a week or so?

* get ready to swap G&H mirrors with LaserOptik mirrors (PR2, PR3, SR2, SR3)

* assembly of black glass beam dumps - one afternoon

* green periscope moving - day or so to make sure greens are resonant in arms, so we have something to align to

* redraw of REFL path?? (rotate 2nd refl mirror to reflect beam to the east, then add mirror to get beam out of IOO chamber viewport), cut hole in BS oplev table's box, install black beam tube

* FI camera mount - make sure it's here, baked

* active TT pitch damping plan

* passive TT weight addition and pitch alignment plan

* camera plan for taking in-chamber photos without touching the tables

* look at layout of ITMY table.  POY pick-off too close to main beam.  Can we move POY pick-off to BS table?

* remount black glass baffles on SOS cages.

 

Stuff to bake:

* Quadrupus cables

* TT base plates

* FI mirror mount for FI camera

* stuff to add to active TTs

* stuff to add to passive TTs

 

  7557   Tue Oct 16 11:54:05 2012 JenneUpdateIOOMC alignment??

The MC won't survive the boosts right now.  Pizza meeting is in a minute, and I won't be back to the lab before ~3:30 because of the seminar / a meeting, so someone else is welcome to try to fix it. Otherwise I'll have a look later on.

I'm leaving the autolocker disabled, so that it won't try any funny business.  WFS are off, so that they don't need to be turned off by the down script.

  7556   Tue Oct 16 11:38:17 2012 JenneUpdateLSCMore PRMI notes from last night

Quote:

How can you lock the PRMI without the REFL beams? c.f. this entry by Kiwamu
Which signals are you using for the locking?

I think the first priority is to find the fringes of the arms and lock them with POX/POY.

As for the POP, make sure the beam is not clipped because the in-vac steering mirrors
have been supposed to be too narrow to accommodate these two beams.

I was using AS55I for PRCL, and AS55Q for MICH.  I snuck that into the last line of an unrelated elog, since I did both things at the same time: see elog 7551.  Kiwamu's measurements (elog 6283) of the PRMI sensing matrix show that the PRCL and MICH signals are almost orthogonal in AS55 (although the optickle simulation doesn't agree with that...)  He was able to lock PRMI with AS55 I&Q (elog 6293), so I thought we should be able to as well.  Locking the PRMI was supposed to help tune the alignment of the PRM, not be the end goal of the night.  Also, we only tried locking PRCL in the "middle right" configuration, not the "lower left" configuration, but we were seeing what looked like recycling flashes only in the "lower left" configuration.

I agree in principle that we should be working on the arms. However, since we can't use the old steer-the-beam-onto-the-cage trick to find the beam, I was hoping that we could steer the beam around and see some light leaking out of the ETM, onto the end table.  However, with the 1% transmission of the ITMs and ~10ppm transmission of the ETMs, there's not a lot of light back there.  I was hoping to align the PRMI so that I get flashes with a gain of 10 if I'm lucky, rather than just the 5% transmission of the PRM.  With the PRMI aligned, I was expecting:

(1W  through Faraday) * (10 PR gain) * (0.5 BS transmission) * (0.01 ITM transmission) * (10ppm ETM transmission) = 0.5uW on the ETM tables during PRCL flashes

I was hoping that things would be well enough aligned that I could just go to the end table, and see the light with a viewer, although as I type this, I realize that if the beam was not on the end table (or even if it was...) any time I move the PZTs, I'd have to completely realign the PRMI in order to see the flashes.  This seems untenable, unless there are no other options.

We then got sidetracked by trying to see the POP beam, and once we saw the POP beam we wanted to put something down so we could find it again.  POP is also small, but not as small as expected at the end:

(1W  through Faraday) * (10 PR gain) * (20ppm PR2 transmission) = 0.2mW on POP during PRCL flashes.

POP was very difficult to see, and we were only able to see it by putting the foil in the beam path, and using a viewer.  I think that we once were able to see it by looking at a card with the viewer, but it's much easier with the foil.  I'd like to find an iris that is shiny (the regular black iris wasn't helpful), to facilitate this alignment.  Since we were just looking at the reflection off of the foil, I have no comment yet about clipping vs. not clipping.  I do think however that the forward-going beam may have been easier to find....when the PRMI alignment drifted, we lost the beam, but I could still see the forward-going beam.  Probably I should switch to that one, since that's the one that was lined up with the in-vac optics. 

Summary:

Ideas are welcome, for how to align the beam to the Yarm (and later to the Xarm), since our old techniques won't work.  Aligning the PRMI was a distraction, although in hopes of getting flashes so we could see some light at the end tables.  I'm going to go see if I can look through a viewport and see the edges of the black glass aperture, which will potentially be a replacement for the steering-on-the-cage technique, but if that doesn't work, I'm running out of ideas.

  7555   Tue Oct 16 02:34:38 2012 KojiUpdateLSCPOP lens placed on POX table, 2 PRMI movies

How can you lock the PRMI without the REFL beams? c.f. this entry by Kiwamu
Which signals are you using for the locking?

I think the first priority is to find the fringes of the arms and lock them with POX/POY.

As for the POP, make sure the beam is not clipped because the in-vac steering mirrors
have been supposed to be too narrow to accommodate these two beams.

  7554   Tue Oct 16 00:33:29 2012 JenneUpdateLSCPOP lens placed on POX table, 2 PRMI movies

[Evan, Jenne]

We aligned the PRMI.  We definitely can lock MICH, but we're not really sure if PRCL is really being locked or not.  I don't think it is.

Anyhow, we found 2 different places on the AS camera that we can align the PRMI.  One (middle, right hand side of the camera), we see the same weird fringing that we've been seeing for a week or two.  The other (lower left side of the camera), we see different fringing, almost reminds me more of back in the day a few months ago when the beam looked like it was expanding on each pass.  As I type, Evan is uploading the movies to youtube.  I *still* don't know how to embed youtube videos on the elog!

Also, we found both forward-going and backward-going POP beams coming out onto the POX table.  We placed the 2" lens in the path of the backwards beam, so that we can find it again.  We can't see it on an IR card, but if we put some foil where we think the beam should be, we can use a viewer to see the spot on the foil.  Poking a hole in the foil made an impromptu iris.

Youtube videos:

Lower left on camera

Middle right on camera

  7553   Tue Oct 16 00:08:26 2012 DenUpdateIOOc1lsc DAC0 now connected to tip-tilt SOS DW boards

Quote:

Tomorrow I'll get new SMA cables to connect the DW/AI outputs to the coil driver boards, and I'll start testing the coil driver outputs. 

 I've found a nice 16 twisted pair cable ~25m long and decided to use it as a port from 1Y3 to clean room cable instead of buying a new long one. I've added a break out board to the coil driver end to monitor outputs.

DSC_4748.JPG

  7552   Mon Oct 15 22:24:45 2012 JenneUpdateComputersLots of new White :(

Evan and I are starting to lock, and there is lots of new, unfortunate white stuff on several different screens.

C1:TIM-PACIFIC_STRING is gone, C1:IFO-STATE (MC state) is gone, C1:LSC-PZT..._requests are gone (all 4 of them), C1:PSL-FSS_FASTSWEEPTEST from the FSS screen is gone (although I'm not sure that that one is newly gone), lots of the WF AA lights on the LSC screen are gone.

Those are the things I find in a few minutes of not really looking around.

EDIT:  IPPOS is also gone, so I can't see how my current alignment relates to old alignments.

  7551   Mon Oct 15 22:16:09 2012 JenneUpdateSUSAll oplev gains turned to 0

Steve has promised to fix up all of the oplevs, but it hasn't happened yet, so I've turned all of the oplev gains to zero, so that when the optics are restored we don't have to quickly click them off.

Oplev values that were changed to zero:

PRM P=0.15, Y=-0.3

SRM P=-2.0, Y=2.0

BS P=0.2, Y=-0.2

ITMY P=2.1, Y=-2.0

ITMX P=1.0, Y=-0.5

ETMX P=-0.2, Y=-0.2

ETMY P=0.5, Y=0.6

Also, PRCL was changed in the LSC input matrix from REFL33I to AS55I, since there is no REFL beam out of the IFO :(

  7550   Mon Oct 15 20:45:58 2012 jamieUpdateIOOc1lsc DAC0 now connected to tip-tilt SOS DW boards

The tip-tile SOS dewhite/AI boards are now connected to the digital system.

20121015_190340.png

I put together a chassis for one of our space DAC -> IDC interface boards (maybe our last?).  A new SCSI cable now runs from DAC0 in the c1lsc IO chassis in 1Y3, to the DAC interface chassis in 1Y2.

Two homemade ribbon cables go directly from the IDC outputs of the interface chassis to the 66 pin connectors on the backplane of the Eurocrate.  They do not go through the cross-connects, cause cross-connects are stupid.  They go to directly to the lower connectors for slots 1 and 3, which are the slots for the SOS DW/AI boards.  I had to custom make these cables, or course, and it was only slightly tricky to get the correct pins to line up.  I should probably document the cable pin outs.

  • cable 0:  IDC0 on interface chassis (DAC channels 0-7) ---> Eurocrate slot 0 (TT1/TT2)
  • cable 1:  IDC1 on interface chassis (DAC channels 8-15)---> Eurocrate slot 2 (TT3/TT4)

As reported in a previous log in this thread, I added control logic to the c1ass front-end model for the tip-tilts.  I extended it to include TT_CONTROL (model part) for TT3 and TT4 as well, so we're now using all channels of DAC0 in c1lsc for TT control.

I tested all channels by stepping through values in EPICS and reading the monitor and SMA outputs of the DW/AI boards.  The channels all line up correctly.  A full 32k count output of a DAC channel results in 10V output of the DW/AI boards.  All channels checked out, with a full +-10V swing on their output with a full +-32k count swing of the DAC outputs.

   We're using SN 1 and 2 of the SOS DW/AI boards (seriously!)

The output channels look ok, and not too noisy.

Tomorrow I'll get new SMA cables to connect the DW/AI outputs to the coil driver boards, and I'll start testing the coil driver outputs.

As a reminder:  https://wiki-40m.ligo.caltech.edu/Suspensions/Tip_Tilts_IO

 

  7549   Mon Oct 15 17:08:01 2012 SteveUpdateVACvacuum is back with some disabled valves

Quote:

Quote:

Apparently all of the ION pump valves (VIPEE, VIPEV, VIPSV, VIPSE) opened, which vented the main volume up to 62 mTorr.  All of the annulus valves (VAVSE, VAVSV, VAVBS, VAVEV, VAVEE) also appeared to be open.  One of the roughing pumps was also turned on.  Other stuff we didn't notice?  Bad. 

 Several of the suspensions were kicked pretty hard (600+ mV on some sensors) as a result of this quick vent wind.  All of the suspensions are damped now, so it doesn't look like we suffered any damage to suspensions.

CLOSE CALL on the vacuum system:

Jamie and I disabled V1, VM2 and VM3 gate valves by disconnecting their 120V solenoid actuator before the swap of the VME crate.

The vacuum controller unexpectedly lost control over the swap as Jamie described it. We were lucky not to do any damage! The ion pumps were cold and clean. We have not used them for years so their outgassing possibly  accumulated to reach ~10-50 Torr

I disconnected_ immobilized and labelled the following 6 valves:  the 4 large ion pump gate valves and VC1,  VC2  of the cryo pump. Note: the valves on the cryo pump stayed closed. It is crucial that a warm cry pump is kept closed!

This will not allow the same thing to happen again and protect the IFO from warm cryo contamination.

The down side of this that the computer can not identify vacuum states any longer.

This vacuum system badly needs an upgrade. I will make a list.

Attachment 1: VMEcrateout.png
VMEcrateout.png
Attachment 2: IMG_1725.JPG
IMG_1725.JPG
Attachment 3: warning.png
warning.png
Attachment 4: shortcomings.jpg
shortcomings.jpg
  7548   Mon Oct 15 14:51:16 2012 JenneUpdateSUSSUS were kicked hard as a result

Quote:

Apparently all of the ION pump valves (VIPEE, VIPEV, VIPSV, VIPSE) opened, which vented the main volume up to 62 mTorr.  All of the annulus valves (VAVSE, VAVSV, VAVBS, VAVEV, VAVEE) also appeared to be open.  One of the roughing pumps was also turned on.  Other stuff we didn't notice?  Bad. 

 Several of the suspensions were kicked pretty hard (600+ mV on some sensors) as a result of this quick vent wind.  All of the suspensions are damped now, so it doesn't look like we suffered any damage to suspensions.

  7547   Mon Oct 15 13:07:51 2012 jamieUpdateVACvacuum VME crate broken, replaced, minor vacuum mayhem ensues

Steve and I managed to access the fuse in the vacuum VME crate, but replacing it did not bring it back up.  We decided to replace the entire crate.

We manually checked that the most important valves, VC1, VM1 and V1, were all closed.  We disconnected their power so that they would automatically close, and we wouldn't have to worry about them accidentally opening when we rebooted the system.

We noted where all the cables were, disconnected everything, and removed the crate.  We noted that one of the values switched when we disconnected one of the IPC cables from a VME card.  We'll note which one it was in a followup post.  We thought that was a little strange, since the VME crate was completely unpowered.

Anyway, we removed the crate, swapped in a spare, replaced all the cards and connections, double checked everything, then powered up the crate.  That's when minor chaos ensued.

When the system came back online after about 20 seconds, we heard a whole bunch of valves switching.  Luckily we were able to get the medm screens back up so that we could see what was going on.

Apparently all of the ION pump valves (VIPEE, VIPEV, VIPSV, VIPSE) opened, which vented the main volume up to 62 mTorr.  All of the annulus valves (VAVSE, VAVSV, VAVBS, VAVEV, VAVEE) also appeared to be open.  One of the roughing pumps was also turned on.  Other stuff we didn't notice?  Bad.

We ran around and manually unplugged all of the ION pump valves, since I couldn't immediately pull up the vacuum control screen.  Once that was done and we could see that the main volume was closed off we went back to figure out what was going on.

We got the medm vacuum control screen back (/cvs/cds/caltech/medm/c0/ve/VacControl_BAK.adl.  really??)  There was a lot of inconsistency between the readback states of the valves and the switch settings.  Toggling the switches seemed to bring things back in line.  At this point it seemed that we had control of the system again.  The epics readings were consistent with what we were seeing in the vacuum rack.

We went through and closed everything that should have been closed.  The line pressure between the big turbo pump TP1 and the rest of the pumps was up at atmosphere, 700 Torr.  We connected the roughing pumps and pumped down the lines so that we could turn the turbos back on.  Once TP2 and TP3 were up to speed, we turned on TP1 and opened V1 to start pumping the main volume back town.   The main volume is at 7e-4 Torr right now.

 

So there are a couple of problems with the vacuum system.

  • Why the hell did valves open when we rebooted the VME crate?  That's very bad.  That should never happen.  If the system is set to come up to an unsafe state that needs to be fixed ASAP.  The ION pump valves should never have opened.  Nor the annulus valves.
  • Why were the switches and the readbacks showing different states?
  • Apparently there is no control of the turbo pumps through MEDM.  This should be fixed.

I connect belledona, the laptop at the vacuum station to the wired network, so that it's connection would be less flaky.

 

  7546   Mon Oct 15 10:14:22 2012 DenUpdatePEMAA board

Quote:

 

 I'll leave accelerometers near ETMX table. It's interesting to measure table motion in the morning when trucks drive by.

 That low frequency effect was due to AA board, now it is gone.

  7545   Mon Oct 15 08:12:45 2012 SteveUpdateVACblank vac monitor

VME crate  blew its fuse and the fuse housing is hard to access.

Attachment 1: blankvac.png
blankvac.png
  7544   Mon Oct 15 08:08:33 2012 steveUpdateSUSPRM damping restored

Quote:

Quote:

Quote:

Quote:

Yuta claims he fixed the PRM oplev by centering it the other day, but no one has left it on and watched it for a long while, to make sure it's okay.  We watched it now for ~2 min, and it was good, but we're leaving the oplevs off anyway for the night.  Tomorrow we should restore PRM (it's currently restored), turn on the oplevs, and let it sit to make sure it doesn't go crazy.

 

 PRM oplev servo was turned on with PITgain 0.5  and YAWgain  -0.7

Note: gain settings were PIT  1.0  and  YAW --0.5   on Jun 1, 2012 that I measured Feb 23, 2012

 It is still oscillating. Gains turned down to zero.

 Earthquake test our suspensions            PRM damping restored.             Oplev servo gains turned to zero.

 The PRM damping restored. Oplev  PIT gain  0.15 and YAW  gain   -0.3 turned to zero.

  7543   Sun Oct 14 20:51:20 2012 DenUpdatePEMAA board

Quote:

 

 But every ~40 min ETMX motion is much higher then ground motion at low frequencies (<5 Hz). I wonder if this a reaction of a table to outside disturbances or accelerometer issue. 

 This could come from AA board, its range is +/- 2.5 V, RMS of the ETMX table motion is a few times higher then ground motion, so ETMX accelerometer signal was corrupted.

acc.png

As this small AA range has already caused problems before, I decided to increase it. I've looked through the board scheme and found that all its differential line receives and output amplifiers have absolute maximum range of 40V. We used KEPKO power supply for this board with a voltage range up to 6 V. So I've replaced it with a BK PRECISION power supply and set it to +/- 15 V. Now AA board range is 7.5 V.

DSC_4747.JPG 

I'll leave accelerometers near ETMX table. It's interesting to measure table motion in the morning when trucks drive by.

  7542   Sun Oct 14 17:26:03 2012 DenUpdatePEMETMX table

Quote:

 I've measured ETMX table motion compared to ground motion using accelerometers. Data and settings in the xml file are at the svn directory 40m_seismic/etmx

 High frequency (>60 Hz) resonances that are present at the ETMX motion spectrum seem to be understandable. Amplification ETMX/GROUND of a factor of 2 at 1 Hz is interesting. I've monitored ACC DQ channels for a few hours and noticed that usually spectrum looks like in the previous elog. But every ~40 min ETMX motion is much higher then ground motion at low frequencies (<5 Hz). I wonder if this a reaction of a table to outside disturbances or accelerometer issue.

 

Attachment 1: etmx.pdf
etmx.pdf etmx.pdf etmx.pdf etmx.pdf
  7541   Sun Oct 14 16:44:20 2012 DenUpdatePEMETMX table

 I've measured ETMX table motion compared to ground motion using accelerometers. Data and settings in the xml file are at the svn directory 40m_seismic/etmx.

DSC_4742.jpg

Attachment 2: 14OCT2012.pdf
14OCT2012.pdf 14OCT2012.pdf 14OCT2012.pdf 14OCT2012.pdf
Attachment 3: 14OCT2012.xml
<?xml version="1.0"?>
<!DOCTYPE LIGO_LW [
<!ELEMENT LIGO_LW ((LIGO_LW|Comment|Param|Time|Table|Array|Stream)*)>
<!ATTLIST LIGO_LW Name CDATA #IMPLIED Type CDATA #IMPLIED>
<!ELEMENT Comment (#PCDATA)>
<!ELEMENT Param (#PCDATA)>
<!ATTLIST Param Name CDATA #IMPLIED Type CDATA #IMPLIED Dim CDATA #IMPLIED
                Unit CDATA #IMPLIED>
<!ELEMENT Table (Comment?,Column*,Stream?)>
<!ATTLIST Table Name CDATA #IMPLIED Type CDATA #IMPLIED>
... 231779 more lines ...
  7540   Sun Oct 14 11:41:42 2012 DenUpdatePEMETMY table

Quote:

 

Accelerometers were installed on the ETMY table and nearby ground to measure amplification of the seismic noise due to the table. During this experiment ground and table motions were measured simultaneously

 I've added xml file with measurement settings and data to 40m svn at directory 40m_seismic/etmy.

DSC_4739.JPG 

Attachment 2: 14OCT2012.pdf
14OCT2012.pdf 14OCT2012.pdf 14OCT2012.pdf 14OCT2012.pdf
Attachment 3: 14OCT2012.xml
<?xml version="1.0"?>
<!DOCTYPE LIGO_LW [
<!ELEMENT LIGO_LW ((LIGO_LW|Comment|Param|Time|Table|Array|Stream)*)>
<!ATTLIST LIGO_LW Name CDATA #IMPLIED Type CDATA #IMPLIED>
<!ELEMENT Comment (#PCDATA)>
<!ELEMENT Param (#PCDATA)>
<!ATTLIST Param Name CDATA #IMPLIED Type CDATA #IMPLIED Dim CDATA #IMPLIED
                Unit CDATA #IMPLIED>
<!ELEMENT Table (Comment?,Column*,Stream?)>
<!ATTLIST Table Name CDATA #IMPLIED Type CDATA #IMPLIED>
... 231759 more lines ...
  7539   Fri Oct 12 22:44:49 2012 DenUpdatePEMETMY table

Quote:

Seismic noise on the ETMY table measured to be a few times higher then on the floor in horizontal direction in the frequency range 50 - 200 Hz. Attached are compared spectrums of X, Y and Z motions.

Accelerometers were installed on the ETMY table and nearby ground to measure amplification of the seismic noise due to the table. During this experiment ground and table motions were measured simultaneously.

DSC_4734.JPG     DSC_4736.JPG

Attachment 1: etmy_x_psd.pdf
etmy_x_psd.pdf
Attachment 2: etmy_y_psd.pdf
etmy_y_psd.pdf
Attachment 3: etmy_z_psd.pdf
etmy_z_psd.pdf
Attachment 4: etmy_coh.pdf
etmy_coh.pdf etmy_coh.pdf etmy_coh.pdf
  7538   Fri Oct 12 16:59:26 2012 steveUpdateSUSTT PITCH damping

Quote:

Quote:

 We can only decide the need of pitch damping when the coils are activated.

Could you tell us why? Are you thinking about induced current damping?

      I  was wrong The instability will be the same when the coils are actuated.

  7537   Fri Oct 12 15:31:03 2012 RijuparnaConfiguration cavitymode scan

 Rijuparna, Manasa

Today I have checked the optical layout of the MC transmission RFPD table and measured the laser powers at different points. Manasa helped me for that. I found the power entering the RF photodiode is 0.394mW while the transmitted power of the cavity is 2.46mW. (I will give the diagram later).

  7536   Fri Oct 12 14:01:33 2012 KojiUpdateSUSTT PITCH damping

Quote:

 We can only decide the need of pitch damping when the coils are activated.

Could you tell us why? Are you thinking about induced current damping?

  7535   Fri Oct 12 10:39:45 2012 SteveUpdateSUSTT PITCH damping

Quote:

Quote:

  [Koji, Steve, Den]

TT alignment is fine, yaw damping is satisfactory, pitch damping is slow. We might want to add magnets to the mirror and attach blades to the frame for pitch edge current damping.

We are moving towards electronics testing.

Atm1,  TT 1.5" high adaptor base will be back from the shop in 10 days.

Atm2,  There is no PITCH damping, YAW edie current damping works well at 0.5 mm gap

Atm3,  Adjustable Al -disc that contains a small magnet is purely designed.

We have to come up with a solution to have damping in PITCH

 We can only decide the need of pitch damping when the coils are activated.

Attachment 1: IMG_1709.JPG
IMG_1709.JPG
  7534   Fri Oct 12 01:56:26 2012 kiwamuUpdateGeneralAlignment situation of interferometer

[Koji / Kiwamu]

 We have realigned the interferometer except the incident beam.

 The REFL beam is not coming out from the chamber and is likely hitting the holder of a mirror in the OMC chamber.

So we need to open the chamber again before trying to lock the recycled interferometers at some point.

 

--- What we did

  •  Ran the MC decenter script to check the spot positions.
    • MC3 YAW gave a - 5mm offset with an error of about the same level.
    • We didn't believe in this dither measurement.
  •  Checked the IP-POS and IP-ANG trends.
    • The trends looked stable over 10 days (with a 24 hours drift).
    • So we decided not to touch the MC suspensions.
  • Tried aligning PRM
  • Found that the beam on the REFL path was a fake beam
    • The position of this beam was not sensitive to the alignment of PRM or ITMs.
    • So certainly this is not the REFL beam.
    • The power of this unknown beam is about 7.8 mW
  • Let the PRM reflection beam go through the Faraday
    • This was done by looking at the hole of the Faraday though a view port of the IOO chamber with an IR viewer.
  • Aligned the rest of the interferometer (not including ETMs)
    • We used the aligned PRM as the alignment reference
    • Aligned ITMY such that the ITMY reflection overlaps with the PRM beam at the AS port.
    • Aligned the BS and SRM such that their associated beam overlap at the AS port
    • Aligned ITMX in the same way.
    • Note that the beam axis, defined by the BS, ITMX  and SRM, was not determined by this process. So we need to align it using the y-arm as a reference at some point.
    • After the alignment, the beam at the AS port still doesn't look clipped. Which is good.

 

---- things to be fixed

   - Align the steering mirrors in the faraday rejected beam path (requires vent)

   - SRM oplev (this is out of the QPD range)

   - ITMX oplev (out of the range too)

  7533   Thu Oct 11 21:26:40 2012 janoschUpdateGeneralPRC phase maps

Just some plots. There is nothing new here except for the fact that I learned how to analyze phase maps myself and how to prepare them for Finesse. In other words, everything is ready for a Finesse simulation.

These phase maps show the raw measurement of ITMY, ITMX and PRC:

ITM03_HR_raw.pngITM04_HR_raw.pngPRM02_HR_raw.png

Subtracting out the tilt from all phase maps, and the curvature from the PRC (I found the fit 121m consistent with previous fits), the one obtains the following residuals that can be used in Finesse (order is again ITMY, ITMX and PRC):

ITM03_HR_untilted.pngITM04_HR_untilted.pngPRM02_HR_untilted_uncurved.png

Attachment 3: PRC_40m.png
PRC_40m.png
  7532   Thu Oct 11 14:40:20 2012 DenUpdateSUSdiagonalization

Quote:

MC1 diagonalization is poor, better then before, but still pitch is seen in pos and yaw. Either smth is malfunctioning or flags touch sensors and do not move freely. On the plot mc1_new black lines - before, red - after rediagonalization.

 I've manually corrected MC1 input matrix by looking at UL, UR, LL, LR transfer functions between each other. This improved pos significantly and slightly yaw.

Attachment 1: mc1.pdf
mc1.pdf
  7531   Thu Oct 11 12:11:23 2012 jamieUpdateIOOc1ass with new DAC0 output has been recompiled/install/restarted

I rebuilt/install/restarted c1ass.  It came up with no problems.  It's now showing DAC0 with no errors.

After lunch I'll test the outputs.

  7530   Thu Oct 11 12:02:15 2012 DenUpdateIOOFSS

FSS SLOW control did not drift during the lock at night with MCL path working and AC coupled.

fss.png

  7529   Thu Oct 11 11:57:40 2012 jamieUpdateCDSall IOP models rebuild, install, restarted to reflect fixed ADC/DAC layouts

Quote:

As Rolf pointed out when he was here yesterday, all of our IOPs are filled with parts for ADCs and DACs that don't actually exist in the system.  This was causing needless module error messages and IOP GDS screens that were full of red indicators.  All the IOP models were identically stuffed with 9 ADC parts, 8 DAC parts, and 4 BO parts, even though none of the actual front end IO chassis had physical configurations even remotely like that.  This was probably not causing any particular malfunctions, but it's not right nonetheless.

I went through each IOP, c1x0{1-5}, and changed them to reflect the actual physical hardware in those systems.  I have committed these changes to the svn, but I haven't rebuilt the models yet.  I'll need to be able to restart all models to test the changes, so I'm going to wait until we have a quiet time, probably next week.

I finally got around to rebuilding, installing, and restarting all the IOP models.  Everything went smoothly.  I had to restart all the models on all the screens, but everything seemed to come back up fine.  We now have many fewer dmesg error messages, and the GDS_TP screens are cleaner and don't have a bunch of needless red.

A frame builder restart was also required, due to name changes in unused (but unfortunately still needed) channels in the IOP.

  7528   Thu Oct 11 11:36:52 2012 SteveUpdateSUSTT pitch adjustment

Tip Tilt pitch adjustment on existing-in vacuum suspension. This can be added by a simple installation of a 1.25" long 2-56 threaded rod with nuts.

Attachment 1: IMG_1702.JPG
IMG_1702.JPG
  7527   Thu Oct 11 11:20:05 2012 janoschUpdateGeneralbeam shape simulation, PRC

I started to create a Finesse model of the PRC cavity. We have the phase maps for the PRC and the two ITMs. I could not find anything for PR2,3 and BS. All files can be found in my SVN folder /janosch/PRC40m. I used the AutoCAD model to determine angles of incidence and distances. These numbers are largely inconsistent with numbers that you can find elsewhere on the 40m wiki, but this certainly depends on what accuracy is required for interferometer alignment and I don't understand anything about alignment.

The phase maps come in a format that needs to be modified before they can be used in Finesse. I have started with this work, but maybe someone else can take over. The phase maps show tilts and the PRC also has the curvature. These have to be subtracted out before the maps can be loaded into Finesse. I asked GariLynn for the code that they use. The Finesse model (MichPRC_40m.kat) does not load the phasemaps yet, and I just wrote some random parameter values for the TEM00 input beam to the PRC. So these Gauss parameters need to be corrected.

I will only go on with this work if Rana tells me that I should do so, otherwise it is on hold until we have a volunteer.

  7526   Thu Oct 11 01:30:11 2012 DenUpdateSUSdiagonalization

Quote:

MC1 diagonalization is poor, better then before, but still pitch is seen in pos and yaw. Either smth is malfunctioning or flags touch sensors and do not move freely. On the plot mc1_new black lines - before, red - after rediagonalization.

 I've actuated on MC1 with UL, UR, LR, LL coils in turn and measured sensor readings. All coils separately work fine from the first look.

On the plot: black - free mirror, blue - UL coil actuation, green - UR, grey - LR, red - LL.

Attachment 1: mc1_coils.pdf
mc1_coils.pdf
  7525   Thu Oct 11 00:28:30 2012 DenUpdateSUSdiagonalization

I've written MC123 input matrixes to the front-end.

MC1 diagonalization is poor, better then before, but still pitch is seen in pos and yaw. Either smth is malfunctioning or flags touch sensors and do not move freely. On the plot mc1_new black lines - before, red - after rediagonalization.

Attachment 1: mc1_new.pdf
mc1_new.pdf
Attachment 2: mc123_new.pdf
mc123_new.pdf
  7524   Thu Oct 11 00:22:58 2012 SUS_DiagonalizerUpdateSUSOptics kicked

Quote:
All suspended optics have been kicked at Wed Oct 10 21:50:39 PDT 2012. Watchdogs will be reengaged in 90 minutes.


New SUS input matrix diagonalization complete.
  7523   Wed Oct 10 21:50:39 2012 SUS_DiagonalizerUpdateSUSOptics kicked
All suspended optics have been kicked at Wed Oct 10 21:50:39 PDT 2012. Watchdogs will be reengaged in 90 minutes.
  7522   Wed Oct 10 20:27:40 2012 DenUpdateIOOMCL, WFS triggers

I've added MCL and WFS stop triggers into C1MCS/SUS model. Threshold value of MC_TRANS can be changed in the text entry located in MC2_POSITION medm screen. I tried 2 cases: trigger either blocks signal before MCL filter bank input or after output. Due to filter history in the 1 case MC2 was still slightly disturbed (C1:SUS-MC2_ULPD_VAR ~= 15) right after unlock. In the second case there was no disturbance as we zero output signal, but then I had to add "clear history" command to the mcup script.

WFS triggers block the signal before ASCPIT/YAW filter bank.

MC2_POS.png

Attachment 2: mcl.pdf
mcl.pdf
  7521   Wed Oct 10 19:22:03 2012 jamieUpdateIOOAdded control for input tip-tilts to c1ass

I have added some control logic and appropriate output DAC channels for the input tip-tilts (TT1 and TT2) to the c1ass model.

The plan is for all the tip-tilt drive electronics to live in a Eurocrate in 1Y2.  They will then interface with a DAC in c1lsc.

c1ass runs on the c1lsc front-end machine, and therefore seemed like an appropriate place for the control logic to go.

I added and interface to DAC0, and a top_named IOO block, to c1ass:

2012-10-10-185707_566x330_scrot.png

The IOO block includes two TT_CONTROL library parts, one for each of TT1 and TT2:

2012-10-10-191013_470x277_scrot.png

This is just a start so that I can start testing the DAC output.

I have not recompiled c1ass yet.  I will do that tomorrow.

ELOG V3.1.3-